Circular No. 7718

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ei IN ANONYMOUS GALAXY P. Martin and W. D. Li, University of California at Berkeley, report

the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope. SN 2001ei is located at α = 23^{h}51^{m}02^{s}.95, δ = +27^{o}10′55′′.6 (equinox 2000.0), which is ∼ 5′′ north of the nucleus of a faint galaxy that is itself located ∼ 56′′ east and 122′′ north of the nucleus of NGC 7768. Magnitudes for SN 2001ei: Aug. 14.4 UT, [19.0; 19.4, 18.7; 24.4, 18.0; 25.4, 17.9; 30.4, 17.8; Sept. 5.4, 18.2; 15.4, 19.1; 17.3, 19.2.

COMETS C/1996 Q2, 2001 Q7, 2001 R2, 2001 R3, 2001 R4, 2001 R5

Measurements by D. Hammer (and by Hammer and D. Biesecker for C/2001 R2) have been reported for six more comets discovered on SOHO website images (cf. IAUC 7694). Comet C/2001 Q7, found by R. Kracht, appeared pointlike with no apparent tail on C2 coronagraph images; this object does not have a Kreutz-type orbit, and Biesecker reports the following V magnitudes for C/2001 Q7: Aug. 21.868 UT, 8.1; 21.893, 8.0; 21.921, 8.4; 21.936, 8.5; 21.977, 8.4; 21.993, 8.6. The remaining comets all appear to be Kreutz sungrazers: C/2001 R2 (found by S. Hönig), C/2001 R4 (found by X. Zhou), and C/2001 R5 (found by D. Johnson), were all visible in both the C2 and C3 instruments, while C/2001 R3 (found by M. Meyer) and C/1996 A2 (found by R. Gorelli) were only detected in C3 images. C/2001 R2 reached peak mag 2.5–3 and showed a tail at least 0^{o}.75 long in C3 images. C/2001 R4 and C/2001 R5 each showed a short tail. Astrometric reductions and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  1996 UT  α_{2000}  δ_{2000}  MPEC
C/1996 A2 Jan. 14.058  19^{h}53.^{m}2  −23^{o}06′  2001-O43
Comet  2001 UT  α_{2000}  δ_{2000}  MPEC
C/2001 Q7 Aug. 21.852  9^{h}57.^{m}3  +12^{o}59′  2001-R36
C/2001 R2 Sept. 3.142  10 14.0  + 5 53  2001-R34
C/2001 R3  5.113  10 36.8  + 6 19  2001-R51
C/2001 R4  8.221  10 50.8  + 4 27  2001-R51
C/2001 R5  10.488  10 52.6  + 4 08  2001-R51
2001 September 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green